# Analysis and Simulation of a Fluid-Heat System in a Thin, Rough Layer in Contact With a Solid Bulk Domain

### v2

@@ 4.2 Homogenized equations
Shifting to the Stokes system, we observe that our setup is somewhat close to fluid flows in porous media. The following steps are therefore similar to the classical derivation of the Darcy equation, see [3, 8]. There is also quite a lot of literature considering thin porous layers, which is even closer related to the present setup, for example, [22, 37, 26]. Nevertheless, to better understand the final homogenized model we briefly present the procedure and point out any differences compared to deriving a Darcy equation.
@@
Let $\varphi\in C_{c}^{\infty}(\Sigma;C_{\#}^{\infty}(\overline{Z}))^{d}$ be a smooth function with compact support in $\Sigma$ and $\varphi(\tilde{x},y)=0$ if $y\in\Gamma\cup\{y_{d}=0\}$. We test the weak form Eq. 3.2 with $\varphi_{\varepsilon}\in H_{0}^{1}(\Omega^{f}_{\varepsilon})^{d}$ given by $\varphi_{\varepsilon}(x)=\varepsilon^{k}\varphi(\tilde{x},\nicefrac{{x}}{{\varepsilon}})$ ($k=1,2$) to get (for almost all $(t\in S)$)
@@
If we set $k=2$ and pass to the limit $\varepsilon\to 0$, we see that
$(p(t,\tilde{x},y),\operatorname{div}_{y}{\varphi}(\tilde{x},y))_{\Sigma\times Z}=0.$
It follows that $p$ is constant in $y$. On the other hand, if we take only functions with $\operatorname{div}_{y}{\varphi}=0$ and $\operatorname{div}_{\tilde{x}}{\left(\int_{Z}\varphi\,\mathrm{d}y\right)}=0$ and set $k=1$, the limit yields
$(\mu(\theta^{f}(t,\tilde{x}))\nabla_{y}u(t,\tilde{x},y),\nabla_{y}\varphi(\tilde{x},y))_{\Sigma\times Z}=0.$ (4.7)
Here, we used Corollary 2. Next, let us introduce
$V=\left\{\varphi\in L^{2}(\Sigma;H^{1}_{\#}(Z))^{d}:\,\begin{array}[]{c}\operatorname{div}_{y}{\varphi}=0,\,\operatorname{div}_{\tilde{x}}{\left(\int_{Z}\varphi\,\mathrm{d}y\right)}=0,\\ \varphi=0\text{ on }\Gamma\cup\{y_{d}=0\}\text{ and }\int_{Z}\varphi\,\mathrm{d}y\cdot n_{\Sigma}=0\text{ on }\partial\Sigma\end{array}\right\}.$
By a density argument Eq. 4.7 also holds for all $\varphi\in V$. By the choice of test functions the pressure $p$ does not appear in (4.7). To recover the pressure, we use the properties of the orthogonal of $V$. With the results from [22, Lemma 5.3 and Remark 5.5.$(i)$], one obtains that
$V^{\perp}=\left\{q=\nabla_{\tilde{x}}q_{0}+\nabla_{y}q_{1}:\,q_{0}\in H^{1}(\Sigma)/\mathbb{R},\,q_{1}\in L^{2}_{0}(\Sigma\times Z)\right\}\subset L^{2}(\Sigma;\mathcal{H}(Z)^{\prime}),$
where $\mathcal{H}(Z)\coloneqq\left\{\varphi\in H^{1}_{\#}(Z)^{d}:\varphi(y)=0\text{ if }y\in\Gamma\cup\{y_{d}=0\}\right\}$. The decomposition into $(q_{0},q_{1})$ is also unique for every $q\in V^{\perp}$. Therefore, there exists a $q=(q_{0},q_{1})\in L^{2}(S;V^{\perp})$ such that $\mu(\theta^{f})\Delta_{y}u=\nabla_{\tilde{x}}q_{0}+\nabla_{y}q_{1}$ (for almost all $t\in S$) understood as an element in $L^{2}(\Sigma;\mathcal{H}(Z)^{\prime})$. Collecting the boundary conditions of $u$ derived in Eq. 4.5 and Proposition 1 as well as the above characterization, we see that
$(\bar{u},u,q_{0},q_{1})\in L^{2}(S;H(\operatorname{div},\Sigma)\times L^{2}(\Sigma;H^{1}(Z))\times H^{1}(\Sigma)/\mathbb{R}\times L^{2}_{0}(\Sigma\times Z))$
@@
[…]
Note that both $u_{\text{motion}}\cdot e_{d}=0$ and Assumption (A5) (which guarantees $\int_{\partial\Sigma}n_{\Sigma}\cdot\int_{0}^{1}u_{BC}=0$) are essential to ensure that the boundary conditions are compatible with the incompressibility requirements. Furthermore, we introduced the new pressure-like functions $q_{0},q_{1}$ that still have no clear connection to the pressure limit $p$ of the microscale model. We first show $p=q_{0}$. To this end, we test the original momentum equation of the micro problem with $\varepsilon\varphi$ for $\varphi\in C_{c}^{\infty}(\Sigma;C_{\#}^{\infty}(\overline{Z}))^{d}$ with $\varphi=0$ on $\Gamma\cup\{y_{d}=0\}$ and $\operatorname{div}_{y}{\varphi}=0$:
$(\mu(\theta^{f})\nabla_{y}u,\nabla_{y}\varphi)_{\Sigma\times Z}-(p,\operatorname{div}_{\tilde{x}}\varphi)_{\Sigma\times Z}=0.$ (4.8)
With a density argument, this also holds for all $\varphi\in L^{2}(\Sigma;\mathcal{H}(Z))$ with $\operatorname{div}_{\tilde{x}}{\left(\int_{Z}\varphi\,\mathrm{d}y\right)}\in L^{2}(\Sigma)$ and $\operatorname{div}_{y}{\varphi}=0$. Since the representation $\mu(\theta^{f})\Delta_{y}u=\nabla_{\tilde{x}}q_{0}+\nabla_{y}q_{1}$ is unique, we deduce from Eq. 4.8 that it needs to hold $p=q_{0}$. We recovered the pressure $p$ for the two-pressure model and therefore replace, in the following, $q_{0}$ by $p$.
By the linearity of the above two-pressure Stokes model, we can further simplify the problem by constructing suitable cell problems. The first set of cell problems are quite standard, see [3] where they are used to characterize the permeability in a porous medium: for $i=1,\dots,d-1$ find $(\xi_{i},\eta_{i})\in H^{1}_{\#}(Z)^{d}\times L^{2}_{0,\#}(Z)$ such that
$-\Delta_{y}\xi_{i}-\nabla_{y}\eta_{i}$ $=e_{i}\quad$ in $\text{in }Z,$
$\operatorname{div}_{y}{(\xi_{i})}$ $=0$ in $\text{in }Z,$ (4.9)
$\xi_{i}$ $=0$ on $\text{on }\Gamma\cup\{y_{d}=0\}.$
@@
$-\Delta_{y}\xi_{0}-\nabla_{y}\eta_{0}$ $=0\quad$ in $\text{in }Z,$
$\operatorname{div}_{y}{(\xi_{0})}$ $=0$ in $\text{in }Z,$ (4.10)
$\xi_{0}$ $=0$ on $\text{on }\Gamma,$
$\xi_{0}$ $=u_{\text{motion}}\quad$ on $\text{on }\{y_{d}=0\}.$
Note, the cell problem (4.2) is a well-defined equation and admits a unique solution, given that $u_{\text{motion}}\cdot e_{d}=0$. We have the following property of the cell solutions:
@@
For the cell solutions $\xi_{i}\in H^{1}_{\#}(Z)^{d}$ ($i=0,\dots,d-1$) of Section 4.2 and Section 4.2, respectively, it holds
@@
Given the solutions of the cell problems (4.2) and (4.2), we obtain for a.a. $(t,\tilde{x},y)\in S\times\Sigma\times Z$ the representations
$u(t,\tilde{x},y)=\xi_{0}(y)-\sum_{i=1}^{d-1}\frac{\xi_{i}(y)}{\mu(\theta^{f}(t,\tilde{x}))}\partial_{x_{i}}p(t,\tilde{x}),\quad q_{1}(t,\tilde{x},y)=\eta_{0}(y)-\sum_{i=1}^{d-1}\frac{\eta_{i}(y)}{\mu(\theta^{f}(t,\tilde{x}))}\partial_{x_{i}}p(t,\tilde{x}).$ (4.11)
This can be checked easily by plugging them into the two-pressure Stokes system. Looking at $\bar{u}=\int_{Z}u\,\mathrm{d}y$, we can also check that with $\operatorname{div}_{\tilde{x}}\bar{u}=0$ and by testing with $\psi\in H^{1}(\Sigma)$ while using Eq. 4.11:
$\begin{split}0&=\left(\operatorname{div}_{\tilde{x}}{\bar{u}},\psi\right)_{\Sigma}=-\left(\bar{u},\nabla_{\tilde{x}}\psi\right)_{\Sigma}+\left(\bar{u}\cdot n_{\Sigma},\psi\right)_{\partial\Sigma}\\ &=-\left(\int_{Z}\xi_{0}\,\mathrm{d}y,\nabla_{\tilde{x}}\psi\right)_{\Sigma}+\sum_{i=1}^{d-1}\left(\frac{\partial_{i}p}{\mu(\theta^{f})}\int_{Z}\xi_{i}\,\mathrm{d}y,\nabla_{\tilde{x}}\psi\right)_{\Sigma}+\left(\left(\int_{0}^{1}u_{BC}(\cdot,s)\,\mathrm{d}s\right)\cdot n_{\Sigma},\psi\right)_{\partial\Sigma}\\ &=-(\bar{\xi}_{0},\nabla_{\tilde{x}}\psi)_{\Sigma}+\left(\frac{1}{\mu(\theta^{f})}K\nabla_{\tilde{x}}p,\nabla_{\tilde{x}}\psi\right)_{\Sigma}+(\bar{u}_{\text{BC}}\cdot n_{\Sigma},\psi)_{\partial\Sigma}.\end{split}$ (4.12)
@@
$\bar{\xi}_{0}\coloneqq\int_{Z}\xi_{0}(y)\,\mathrm{d}y,\quad K_{ij}\coloneqq\int_{Z}\xi_{j}(y)\cdot e_{i}\,\mathrm{d}y\quad\text{and}\quad\bar{u}_{\text{BC}}\coloneqq\int_{0}^{1}u_{BC}(\cdot,s)\,\mathrm{d}s.$ (4.13)
Note, Eq. 4.12 is similar to a Darcy equation, but with an additional term induced by the Dirichlet condition at $y_{d}=0$. The following Theorem 3 collects all equations and shows the complete model.
Theorem 3 (Homogenized model).
Let the Assumptions (A1)-(A7) be fulfilled and let $(\theta^{s},\theta^{f},u,p)$ be the limit functions defined in Proposition 1. Set $\bar{u}(t,\tilde{x})=\int_{Z}u(t,\tilde{x},y)\,\mathrm{d}y$. Then $(\theta^{s},\theta^{f},\bar{u},p)$ are weak solutions of the following problem: inside $\Omega$ it holds
[…]
where $\bar{f}(t,\tilde{x})=\int_{\Gamma}f(t,\tilde{x},y)\,\mathrm{d}\sigma_{y}$. The fluid temperature $\theta^{f}$ fulfills
[…]
with the effective conductivity $\tilde{\kappa}\in\mathbb{R}^{d\times d}$ given by the cell solutions $\{\omega_{i}\}_{i=1}^{d-1}$ of Eq. 4.4, $\omega_{d}=0$ and
$\tilde{\kappa}_{ij}=\int_{Z}\kappa^{f}(\delta_{ij}+\nabla_{y}\omega_{i}\cdot e_{j})\,\mathrm{d}y.$
The fluid movement in the layer is governed by
[…]
here the permeability $K\in\mathbb{R}^{d\times d}$ is given by the solutions $\{\xi_{i}\}_{i=1}^{d-1}$ of the cell problems Section 4.2 via
$K_{ij}=\int_{Z}\xi_{j}\cdot e_{i}\,\mathrm{d}y$
where we set $\xi_{d}=0$ and $\bar{\xi}_{0}=\int_{Z}\xi_{0}(y)\,\mathrm{d}y$, with $\xi_{0}$ given by Section 4.2.
Proof.
That $(\theta^{s},\theta^{f},\bar{u},p)$ fulfill Eq. 4.14 can be derived with the concept of two-scale convergence, as explained and partially carried out above. ∎
Remark 4.
Note, that by Proposition 2 and the equality Eq. 4.14i we obtain that $\bar{u}\cdot e_{d}=0$. This is reasonable since we remain with an averaged interface velocity in the effective model and a nonzero flow in direction $d$ would imply a flow into (or out of) the solid domain. We recovered, ignoring the temperature-dependent viscosity, the same effective equation as in [12], where the needed assumptions for convergence are reduced, thanks to the two-scale concept.
Additionally, in 2D the effective problem for the fluid simplifies further. By Eq. 4.14j it holds $\operatorname{div}_{\tilde{x}}(\bar{u})=\partial_{x_{1}}\bar{u}_{1}=0$ and the velocity is a constant solely given by the boundary function $u_{BC}$ in Eq. 4.14l. This linearizes the problem since the temperature equation is independent of the pressure $p$.
The existence of a solution of model (4.14) is ensured by the existence of the original problem and the two-scale limits. Similar to the micro model, for a unique solution, we need additional regularity assumptions.
Theorem 4 (Uniqueness of the homogenized model).
If there is a solution of the homogenized problem where $\nabla_{\tilde{x}}\theta\in L^{\infty}(S;L^{2+\delta}(\Sigma))$ and $\nabla_{\tilde{x}}p\in L^{\infty}(S;L^{2+\delta}(\Sigma))$ for some $\delta>0$, then this solution is the only.
Proof.
Assume that $(\theta^{f}_{i},\theta^{s}_{i},\bar{u}_{i},p_{i})$, $i=1,2$, are two sets of solutions of the homogenization limit given by System (4.14), i.e., they solve the variational problems given via Eqs. 4.6 and 4.12. In the following, we will denote their differences by $(\bar{\theta}^{f},\bar{\theta}^{s},\bar{\bar{u}},\bar{p})$. W.l.o.g. we assume that $(\theta_{2}^{f},p_{2})$ is the solution enjoying the higher gradient integrability.
From Eq. 4.11 we obtain for $\bar{\bar{u}}$, that
$\|\bar{\bar{u}}\|_{L^{2}(S\times\Sigma)}^{2}\leq C\left(\nabla_{\tilde{x}}\bar{p},\bar{\bar{u}}\right)_{S\times\Sigma}+\left(\left[\frac{1}{\mu(\theta^{f}_{1})}-\frac{1}{\mu(\theta^{f}_{2})}\right]\nabla_{\tilde{x}}p_{2},\bar{\bar{u}}\right)_{S\times\Sigma}$
$\|\bar{\bar{u}}\|_{L^{2}(S\times\Sigma)}\leq C\left(\|\nabla_{\tilde{x}}\bar{p}\|_{L^{2}(S\times\Sigma)}+\|\nabla_{\tilde{x}}p_{2}\|_{L^{\infty}(S;L^{2+\delta}(\Omega))}\|\bar{\theta}^{f}\|_{L^{2}(S;L^{2+\frac{2}{\delta}}(\Sigma))}\right).$
Testing Eq. 4.12 with $\psi=\bar{p}$, we get:
$c\|\nabla_{\tilde{x}}\bar{p}\|_{L^{2}(S\times\Sigma)}\leq C\|\nabla_{\tilde{x}}p_{2}\|_{L^{\infty}(S;L^{2+\delta}(\Omega))}\|\bar{\theta}^{f}\|_{L^{2}(S;L^{2+\frac{2}{\delta}}(\Sigma))}$
$\|\bar{\bar{u}}\|_{L^{2}(S\times\Sigma)}+\|\nabla_{\tilde{x}}\bar{p}\|_{L^{2}(S\times\Sigma)}\leq C\|\bar{\theta}^{f}\|_{L^{2}(S;L^{2+\frac{2}{\delta}}(\Sigma))}.$ (4.15)
We now look at the weak forms for the heat system (Eq. 4.6) and take their difference:
$(\partial_{t}\bar{\theta}^{s},\varphi)_{S\times\Omega}+|Z|(\partial_{t}\bar{\theta}^{f},\psi)_{S\times\Sigma}+\kappa^{s}(\nabla\bar{\theta}^{s},\nabla\varphi)_{S\times\Omega}+(\tilde{\kappa}\nabla_{\tilde{x}}\bar{\theta}^{f},\nabla\psi)_{S\times\Sigma}\\ +(\bar{u}_{1}\cdot\nabla_{\tilde{x}}\bar{\theta}^{f},\psi)_{S\times\Sigma}+\alpha|\Gamma|(\bar{\theta}^{s}-\bar{\theta}^{f},\phi-\psi)_{S\times\Sigma}=-(\bar{\bar{u}}\cdot\nabla_{\tilde{x}}\theta^{f}_{2},\psi)_{S\times\Sigma}$
for all $(\varphi,\psi)\in L^{2}(S;H^{1}(\Omega)\times H^{1}(\Sigma))$. For the advection term on the left-hand side, we test with $(\bar{\theta}^{s},\bar{\theta}^{f})\in L^{2}(S;H^{1}(\Omega)\times H^{1}(\Sigma))$ and exploit that $\bar{u}_{1}$ is divergence-free (cf. Eq. 3.5):
The advective term on the right-hand side can be handled via the additional regularity of $\nabla_{\tilde{x}}\theta^{f}_{2}$. With the same line of argument as in Theorem 2 and using the Sobolev embedding $H^{1}\subset L^{2+\frac{2}{\delta}}$ for 2D domains, uniqueness now follows. ∎

### v3

@@ 4.2 Homogenization limit
Shifting to the Stokes system, we observe that our setup is somewhat close to fluid flows in porous media. The following steps are therefore similar to the classical derivation of the Darcy equation, see [40, 51]. There is also extensive literature considering thin porous layers, which is even closer to the present setup, for example, [35, 52, 49]. Nevertheless, to better understand the final homogenized model, we briefly present the procedure and point out any differences compared to deriving a Darcy equation.
We start by recalling that the $\varepsilon$-solutions $(u_{\varepsilon},p_{\varepsilon})\in V_{\varepsilon}\times Q_{\varepsilon}$ satisfy
$(\mu(\theta_{\varepsilon}^{f})\nabla u_{\varepsilon},\nabla\varphi)_{\Omega^{f}_{\varepsilon}}-(p_{\varepsilon},\operatorname{div}{\varphi})_{\Omega^{f}_{\varepsilon}}=0$ (34)
@@
for all $\varphi\in H_{0}^{1}(\Omega^{f}_{\varepsilon})$ and almost all $t\in S$. Let $\varphi\in C_{c}^{\infty}(\Sigma;C_{\#}^{\infty}(Z))^{d}$ with $\varphi=0$ on $\Sigma\times\Gamma_{d}$. We test the weak form Eq. 34 with $\varphi_{\varepsilon}\in H_{0}^{1}(\Omega^{f}_{\varepsilon})^{d}$ given by $\varphi_{\varepsilon}(x)=\varepsilon^{k}\varphi(\tilde{x},\nicefrac{{x}}{{\varepsilon}})$ and get (for almost all $(t\in S)$)
@@
We set $k=2$ and pass to the limit $\varepsilon\to 0$ with $\varepsilon^{2}p_{\varepsilon}\stackrel{{\scriptstyle 2}}{{\rightharpoonup}}\chi_{Z}p$ (see Proposition 3.(iv)(iv) and note that an additional $\varepsilon^{-1}$ is consumed via the two-scale limit, see Definition 1) to get
$(p(t,\tilde{x},y),\operatorname{div}_{y}{\varphi}(\tilde{x},y))_{\Sigma\times Z}=0$
which, as this holds for all $\varphi\in C_{c}^{\infty}(\Sigma;C_{\#}^{\infty}(Z))^{d}$ with $\varphi=0$ on $\Sigma\times\Gamma_{d}$, implies that $p$ is constant in $y$.
Next, we set $k=1$ and restrict testing to functions for which, additionally, $\operatorname{div}_{y}{\varphi}=0$ holds. The corresponding limit yields
$(\mu(\theta^{f})\nabla_{y}u,\nabla_{y}\varphi)_{\Sigma\times Z}-(p,\operatorname{div}_{\tilde{x}}\varphi)_{\Sigma\times Z}=0$ (35)
where we used $\mu(\theta_{\varepsilon}^{f})\stackrel{{\scriptstyle 2}}{{\rightharpoonup}}\chi_{Z}\mu(\theta_{\varepsilon}^{f})$ (Corollary 2), $\varepsilon\nabla u_{\varepsilon}\stackrel{{\scriptstyle 2}}{{\rightharpoonup}}\chi_{Z}\nabla_{y}u$ (Proposition 3.(iii)(iii)), and $\varepsilon^{2}p_{\varepsilon}\stackrel{{\scriptstyle 2}}{{\rightharpoonup}}\chi_{Z}p$ (Proposition 3.(iv)(iv)). Equation 35 must, by density, also hold for all $\varphi\in L^{2}(\Sigma;H^{1}_{\#,\Gamma_{d}}(Z))$ with $\operatorname{div}_{\tilde{x}}{\left(\int_{Z}\varphi\,\mathrm{d}y\right)}\in L^{2}(\Sigma)$.
Next, let us introduce
$V=\left\{\varphi\in L^{2}(\Sigma;H^{1}_{\#}(Z))^{d}:\,\begin{array}[]{c}\operatorname{div}_{y}{\varphi}=0,\,\operatorname{div}_{\tilde{x}}{\left(\int_{Z}\varphi\,\mathrm{d}y\right)}=0,\\ \varphi=0\text{ on }\Gamma_{d}\text{ and }\int_{Z}\varphi\,\mathrm{d}y\cdot n_{\Sigma}=0\text{ on }\partial\Sigma\end{array}\right\}.$
Naturally, Eq. 35 also holds for every $\varphi\in V$. With the results from [35, Lemma 5.3 and Remark 5.5.(i)(i)], the orthogonal of $V$ is characterized by
$V^{\perp}=\left\{q=\nabla_{\tilde{x}}q_{0}+\nabla_{y}q_{1}:\,q_{0}\in H^{1}(\Sigma)/\mathbb{R},\,q_{1}\in L^{2}_{0}(\Sigma\times Z)\right\}\subset L^{2}(\Sigma;H^{1}_{\#,\Gamma_{d}}(Z)^{\prime})^{d}.$
This decomposition into $(q_{0},q_{1})$ is also unique for every $q\in V^{\perp}$. Therefore, there exists a $q=(q_{0},q_{1})\in L^{2}(S;V^{\perp})$ such that $\mu(\theta^{f})\Delta_{y}u=\nabla_{\tilde{x}}q_{0}+\nabla_{y}q_{1}$ (for almost all $t\in S$) understood as an element in $L^{2}(\Sigma;H^{1}_{\#,\Gamma_{d}}(Z)^{\prime})^{d}$. With Eq. 35 it follows that $q_{0}=p$ almost everywhere up to a constant.
Collecting the boundary conditions of $u$ derived in Eq. 29 and Proposition 3 as well as the above characterization, we see that
$(\bar{u},u,p,q_{1})\in L^{2}(S;H^{1}(\Sigma)\times L^{2}(\Sigma;H^{1}(Z))\times H^{1}(\Sigma)/\mathbb{R}\times L^{2}_{0}(\Sigma\times Z))$
@@
$-\mu(\theta^{f})\Delta_{y}u+\nabla_{y}q_{1}$ $=-\nabla_{\tilde{x}}p\quad$ $\text{ in }S\times\Sigma\times Z,$ (36a)
$\operatorname{div}_{y}{u}$ $=0$ $\text{ in }S\times\Sigma\times Z,$ (36b)
$u$ $=0$ $\text{ on }S\times\Sigma\times\Gamma,$ (36c)
$u$ $=u_{\text{motion}}$ $\text{ on }S\times\Sigma\times\{y_{d}=0\},$ (36d)
$\bar{u}$ $=\int_{Z}u\,\mathrm{d}y$ $\text{ on }S\times\Sigma,$ (36e)
$\operatorname{div}_{\tilde{x}}\bar{u}$ $=0$ $\text{ in }S\times\Sigma,$ (36f)
$\bar{u}(t,\tilde{x})\cdot n_{\Sigma}$ $=\int_{0}^{1}u_{BC}(\tilde{x},s)\,\mathrm{d}s\cdot n_{\Sigma}\quad$ $\text{ on }S\times\partial\Sigma.$ (36g)
Note that both $u_{\text{motion}}\cdot e_{d}=0$ and Assumption (A5) (which guarantees $\int_{\partial\Sigma}n_{\Sigma}\cdot\int_{0}^{1}u_{BC}=0$) are essential to ensure that the boundary conditions are compatible with the incompressibility requirements.
We can further simplify the two-pressure Stokes-problem by constructing suitable cell problems. The first set of cell problems are quite standard, see [40] where they are used to characterize the permeability in a porous medium: for $i=1,\dots,d-1$, find $(\xi_{i},\eta_{i})\in H^{1}_{\#}(Z)^{d}\times L^{2}_{0,\#}(Z)$ such that
$-\Delta_{y}\xi_{i}-\nabla_{y}\eta_{i}$ $=e_{i}\quad$ $\text{in }Z,$
$\operatorname{div}_{y}{(\xi_{i})}$ $=0$ $\text{in }Z,$ (37)
$\xi_{i}$ $=0$ $\text{on }\Gamma_{d}.$
@@
$-\Delta_{y}\xi_{0}-\nabla_{y}\eta_{0}$ $=0\quad$ $\text{in }Z,$
$\operatorname{div}_{y}{(\xi_{0})}$ $=0$ $\text{in }Z,$ (38)
$\xi_{0}$ $=0$ $\text{on }\Gamma,$
$\xi_{0}$ $=u_{\text{motion}}\quad$ $\text{on }\{y_{d}=0\}.$
Note, the cell problem (4.2.2) is a well-defined equation and admits a unique solution, given that $u_{\text{motion}}\cdot e_{d}=0$. We have the following property of the cell solutions:
@@
For the cell solutions $\xi_{i}\in H^{1}_{\#}(Z)^{d}$ ($i=0,\dots,d-1$) of Section 4.2.2 and Section 4.2.2, respectively, it holds
@@
Given the solutions of the cell problems (4.2.2) and (4.2.2), we obtain for a.a. $(t,\tilde{x},y)\in S\times\Sigma\times Z$ the representations
[…]
This can be checked easily by plugging them into the two-pressure Stokes system. Looking at $\bar{u}=\int_{Z}u\,\mathrm{d}y$, we can also check that with $\operatorname{div}_{\tilde{x}}\bar{u}=0$ and by testing Eq. 36a with $\psi\in H^{1}(\Sigma)$ while using Eq. 39:
$\begin{split}0&=\left(\operatorname{div}_{\tilde{x}}{\bar{u}},\psi\right)_{\Sigma}=-\left(\bar{u},\nabla_{\tilde{x}}\psi\right)_{\Sigma}+\left(\bar{u}\cdot n_{\Sigma},\psi\right)_{\partial\Sigma}\\ &=-\left(\int_{Z}\xi_{0}\,\mathrm{d}y,\nabla_{\tilde{x}}\psi\right)_{\Sigma}+\sum_{i=1}^{d-1}\left(\frac{\partial_{i}p}{\mu(\theta^{f})}\int_{Z}\xi_{i}\,\mathrm{d}y,\nabla_{\tilde{x}}\psi\right)_{\Sigma}+\left(\left(\int_{0}^{1}u_{BC}(\cdot,s)\,\mathrm{d}s\right)\cdot n_{\Sigma},\psi\right)_{\partial\Sigma}\\ &=-(\bar{\xi}_{0},\nabla_{\tilde{x}}\psi)_{\Sigma}+\left(\frac{1}{\mu(\theta^{f})}K\nabla_{\tilde{x}}p,\nabla_{\tilde{x}}\psi\right)_{\Sigma}+(\bar{u}_{\text{BC}}\cdot n_{\Sigma},\psi)_{\partial\Sigma}.\end{split}$ (40)
@@
$\bar{\xi}_{0}\coloneqq\int_{Z}\xi_{0}(y)\,\mathrm{d}y,\quad K_{ij}\coloneqq\int_{Z}\xi_{j}(y)\cdot e_{i}\,\mathrm{d}y\quad\text{and}\quad\bar{u}_{\text{BC}}\coloneqq\int_{0}^{1}u_{BC}(\cdot,s)\,\mathrm{d}s.$ (41)
Note, Eq. 40 is similar to a Darcy equation, but with an additional term induced by the Dirichlet condition at $y_{d}=0$. This is precisely the weak form of the Darcy-limit given in Theorem 4.